OpenStudy (anonymous):

How much money invested at 5% compounded continuously for 3 years will yield $820? helpp!

OpenStudy (anonymous):

let the principal be x then u have 820 =x(1+.05)^3 x= 820/(1.05*1.05*1.05)=708.35

OpenStudy (anonymous):

thanks!!!!!! so much:) can u help me with one more?

OpenStudy (anonymous):

write the equation in logarithmic form. 47= 16384

OpenStudy (anonymous):

4^7**

OpenStudy (anonymous):

if its 4^7= 16384 log 16384 to base 4 =7